THE INFRARED SPECTRA OF H_{2}S^{+} AND D_{2}S^{+} TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

When a very dilute mixture of H_{2}S or D_{2}S in neon is codeposited at 4.3 K with a beam of neon atoms that have been excited in a microwave discharge, the infrared spectrum of the resulting solid includes absorptions contributed by the corresponding cations. These assignments are supported both by earlier gas-phase observations and by the results of density functional calculations. The possible assignment of other, less prominent product absorptions to such species as the H_{2}S anion and SH^{+} will also be considered.